OpenStudy (anonymous):

A long jumper leaves the ground at an angle of 20.0° above the horizontal and at a speed of 11.0 m/s. 1. How far does he jump in the horizontal direction? 2. What is the maximum height reached? 3. How far does he jump in the horizontal direction if the jumper leaves the ground at angle 25°, 26°, 27°, 28°, and 29°? 4. What is the maximum height reached f the jumper leaves the ground at angle 25°, 26°, 27°, 28°, and 29°? 5. Plot several graphs for the following information: a. Distance vs angle (Question 3) b. Height vs angle (Que

OpenStudy (anonymous):

Horizontal distance = U^2sin2(theta) / g Maximum height= U^2sin^2(theta) /2g